# Quantum gravitational corrections to Reissner-Nordström black hole thermodynamics and their implications for the weak gravity conjecture

### v3

@@ III.1 The non-extremal RN black hole and the shift of charge-to-mass ratio
The WGC is mainly concerned with the fate of extremal RN black holes. However, we can still explore the quantum gravity effects on non-extremal RN black holes to see how the charge-to-mass ratio changes Solomon:2020pja . In this regard, we find that
$\left.\frac{\partial M}{\partial\epsilon}\right|_{T,q}=-\frac{\mathcal{N}}{300\pi}q^{4}\left(2\bar{r}_{h}^{2}-q^{2}\right)\frac{\ln\left(\bar{r}_{h}^{2}\mu_{*}^{2}\right)}{\bar{r}_{h}^{5}\left(\bar{r}_{h}^{2}-3q^{2}\right)},$ (12)
where $\mathcal{N}$ is defined below in eq.(16). We always have $\mathcal{N}>0$ and $\left(2\bar{r}_{h}^{2}-q^{2}\right)>0$. Then the singularity at $\bar{r}_{h}=\sqrt{3}q$ indicates a critical point that coincides with the boundary between the thermodynamically stable and unstable RN black holes, as determined by the sign of the heat capacity $C_{q}$. Thus, for a thermodynamically stable black hole, the value of $q/M$ increases, while for an unstable one, it decreases. This implies a correlation between the thermodynamic stability and the direction of the shift of charge-to-mass ratio.
III.2 Extermal RN black hole and the WGC
Finally, we are in a position to directly examine the WGC. By taking the extremal limit $T\rightarrow 0$, which is equivalent to $\bar{r}_{h}\rightarrow q$, we get
@@ III.2 Extermal RN black hole and the WGC
Here the parameter $\bar{r}_{h}$ is equal to $q$. In these equations, we have substituted the values of the Wilson coefficients $\beta$, and $\gamma$ from eqs.(4) and (5). As a consistency check, eq.(14) reproduces the logarithmic correction to black hole entropy obtained from an independent approach by Sen in Sen:2012kpz , up to an extra term $-3\ln A$ coming from the zero-modes related to the rotational isometries of an extremal black hole Bhattacharyya:2012wz .
From the expression of $M_{ext}$, we can straightforwardly obtain the result
@@
Interestingly, the coefficients conspire to produce a definite sign. Since $n_{s}$, $n_{f}$ and $n_{v}$ represent the numbers of particles in the spectrum, it is certain that $\mathcal{N}>0$. Consequently, the loop effects of quantum gravity can reduce the mass of an extremal charged black hole at a fixed charge $q$, which was pointed out in Refs.Charles:2019qqt ; Arkani-Hamed:2021ajd .
Though the coefficient $\mathcal{N}$ is independent of the UV theory of quantum gravity, the characteristic energy scale $\mu_{*}$ is determined by the UV theory Calmet:2024neu . Thus, the presence of $\ln\mu_{*}$ in eq.(15) indicates the UV sensitivity of our result. However, the model-dependent value of $\mu_{*}$ does not affect our conclusion for the WGC because we are working within the regime $r>1/\mu_{*}$, where the effective field-theoretical description is valid. The physics near the scale $r\sim 1/\mu_{*}$ would be described by a complete theory of quantum gravity and is not relevant to our current analysis.
Computing $\frac{\partial M_{ext}}{\partial\epsilon}$ is straightforward. However, we can also derive eq.(15) from other formulas, which offer certain advantages. There is a crucial formula, proposed and proven in Goon:2019faz , given by
@@
In fact, from the extended first law $dM=TdS+\Phi dq+U_{\epsilon}d\epsilon$, it can be easily observed that both sides of eq.(17) are equal to $\lim\limits_{T\rightarrow 0}U_{\epsilon}$. This amounts to the evaluation of $\lim\limits_{T\rightarrow 0}(\frac{\partial F}{\partial\epsilon})_{T,q}$, where $F=F_{RN}(T,q)-\epsilon\int d^{3}x\sqrt{g}L_{\text{hd}}$ Xiao:2023two . Therefore, $\frac{\partial M_{ext}}{\partial\epsilon}$ can be evaluated as the spatial integral $-\int d^{3}x\sqrt{g}L_{\text{hd}}$ for an extremal RN black hole Cheung:2018cwt ; Cheung:2019cwi . This provides a rapid way to check whether the WGC is satisfied for a given higher-derivative term $\epsilon L_{\text{hd}}$ without delving into the analysis of the perturbed black hole thermodynamics.
@@
According to the formula (17), another way is to evaluate the shift in entropy due to switching on the quantum gravity effects. This formula is significant as it links the mass decrease required by the WGC to the entropy increase. Notably, our analysis further establishes a novel connection between the required super-extremality and the logarithmic entropy corrections, i.e., both of eqs.(13) and (14) are generated by the same quantum gravity effects. Notice that there is a caveat here. The derivative $T\left(\frac{\partial S}{\partial\epsilon}\right)_{M,q}$ is evaluated at a fixed $M$ rather than at a fixed $T$. Since turning on $\epsilon$ while holding $M$ constant will shift the temperature away from its extremal value, we should use eq.(10) rather than the more familiar eq.(14) to do the calculation. This explains why previously no clues regarding the WGC have been derived from the logarithmic entropy corrections, despite such entropy corrections being known in the literature for many years.
IV A comparison of effective actions
As we mentioned, in the extremal limit, our results regarding super-extremality align with those in Refs.Charles:2019qqt ; Arkani-Hamed:2021ajd . So it is meaningful to conduct a comparison between the effective actions in our study and those in previous works.
@@ IV A comparison of effective actions
where $C_{\mu\nu\alpha\beta}$ represents the Weyl tensor, and the second term exhibits a structure similar to the Gauss-Bonnet term. Given that the RN black hole is Ricci-flat, two independent bases of the higher-derivative terms are sufficient in this context. The combinations of coefficients $\lambda_{1}=\frac{1}{2}(\beta+4\gamma)$ and $\lambda_{2}=-\frac{1}{2}(\beta+2\gamma)$ yield the values
@@
Interestingly, from eqs.(14) and (16), we can observe that $\lambda_{1}$ governs the behavior of super-extremality, while $\lambda_{2}$ controls the logarithmic corrections to the entropy of an extremal RN black hole.
In comparison, the higher-derivative part of the effective action in Refs.Charles:2019qqt ; Arkani-Hamed:2021ajd is often expressed in terms of four-photon interactions $F^{4}$. However, through an appropriate field redefinition, it can also be transformed into the form
@@
Apparently, the two effective actions (18) and (21) seem nearly identical. This is not curious since they ultimately describe the same physical phenomena of loop effects. Nevertheless, in eq.(21), $c_{2}$ appears to be a constant, which may lead to potential confusions. We remind that $c_{2}$ contains the black hole parameter. Therefore, $c_{2}$ needs to be uplifted as an operator similar to the form in eq.(18), when discussing any issues involving back-reactions to the black hole metric. Since the present calculations are carried out based on the unperturbed RN black hole, this does not cause differences in obtaining the results. But it can easily give rise to conceptual problems when interpreting these results.
@@
Firstly, it might be tempting to consider the second term of eq.(21) as a total-derivative term and completely overlook it during the analysis Arkani-Hamed:2021ajd ; Bittar:2024xuc . If doing this, one can still obtain the required super-extremality of WGC but would not find the other side of the story, namely, the logarithmic corrections to the black hole entropy. Secondly, one might explain the absence of $c_{2}$ (equivaletnly $\lambda_{2}$) in eq.(13) by arguing that the second term of eq.(21) is topological and does not affect the equations of motion Charles:2019qqt . However, since the entropy contains a term $\lambda_{2}\ln(\bar{r}_{h}^{2}\mu_{*}^{2})$ for both the non-extremal and extremal cases, the fact is that it is essential for the black hole mass to have a corresponding term $\sim\lambda_{2}\frac{1}{\bar{r}_{h}}$ to maintain the balance; otherwise, the thermodynamic law $dM=TdS+\Phi dq$ cannot hold. This term is not in a logarithmic form, so we did not write it explicitly in eqs.(9) and (13), but it must be there when examining the complete formula. This indicates that the analysis should be more meticulous than initially anticipated.
@@
We now provide a consistent explanation from the perspective of our effective action (18) for why $\lambda_{2}$ appears in the logarithmic term of black hole entropy but not in that of the mass. The second term of eq.(18) can be decomposed into two parts: $\lambda_{2}\ln{\frac{\Box}{\mu_{*}^{2}}}=\lambda_{2}\ln{\Box}-\lambda_{2}\ln{(\mu_{*}^{2})}$. The terms proportional to $\lambda_{2}\ln(\mu_{*}^{2})$ indeed form the Gauss-Bonnet term. It is well-known that a Gauss-Bonnet term can cause a constant shift in the entropy, so it results in an entropy correction of the form $\sim\lambda_{2}\ln(\mu_{*}^{2})$. Due to a matching mechanism between $\ln(\bar{r}_{h}^{2})$ and $\ln(\mu_{*}^{2})$, the entropy actually receives a correction of the form $\sim\lambda_{2}\ln(\bar{r}_{h}^{2}\mu_{*}^{2})$, where $\ln(\bar{r}_{h}^{2})$ stems from the contribution of $\ln\square$. Regarding the expression of mass, the terms proportional to $\lambda_{2}\ln(\mu_{*}^{2})$ are topological and thus do not contribute to its corrections. Then the matching mechanism also prevents the presence of the term $\lambda_{2}\ln{(r_{h}^{2})}$. However, the terms associated with $\lambda_{2}\ln\Box$ is not truly a Gauss-Bonnet term, nothing can preclude them from modifying the metric and the mass formula, only if the correction proportional to $\lambda_{2}$ is not a logarithmic term. Such corrections happen to ensure that the thermodynamic law $dM=TdS+\Phi dq$ holds. See Ref.Xiao:2021zly for a detailed analysis of the variation $\delta\ln\Box$ and its impact on the field equations and black hole solutions.
V Concluding remarks
In this paper, we have investigated the effective action obtained from integrating out massless particles at the one-loop level. Our results confirm that the loop effects of quantum gravity can naturally induce super-extremality ($q>M$) in charged black holes Charles:2019qqt ; Arkani-Hamed:2021ajd . It suggests that quantum gravity plays a crucial role in facilitating the decay of extremal black holes into more fundamental components. These effects thus have far-reaching implications for our understanding of the ultimate evolution of black holes and the universe. We emphasize that the results are model-independent, as they depend solely on the low-energy particle spectrum of nature and are insensitive to the UV details of quantum gravity theories.
We have provided the thermodynamic analysis for both extremal and non-extremal RN black holes, which has interesting implications that warrant further investigation. For extremal RN black holes, our work unexpectedly reveals a connection between the super-extremality of WGC and the logarithmic entropy corrections, which were previously regarded as separate research directions. For non-extremal RN black holes, we establish a relationship between the shift in the charge-to-mass ratio and the thermodynamic stability properties of the black holes. A general discussion on the correlation between thermodynamic stability and super-extremality can be found in Cheung:2018cwt , where it was formalized based on tree-level higher-derivative terms. Our results suggest that this correlation can also be extended to the one-loop effects and non-extremal cases.
Acknowledgments
We thank Grant Remmen, Nima Arkani-Hamed, and Yu-Tin Huang for informing us of their earlier work on loop effects of massless particles and for the constructive communications after our work was posted on arXiv. XY would like to thank Yu Tian and Hongbao Zhang for useful discussions. YX is also thankful to the Higgs Centre for Theoretical Physics at the University of Edinburgh for providing research facilities and hospitality during the visit. This work was supported in part by China Scholarship Council with Grant No. 202408130101, the Hebei Natural Science Foundation with Grant No. A2024201012, and the National Natural Science Foundation of China with Grant No. 12475048.

### v4

@@ 3.1 The non-extremal RN black hole and the shift of charge-to-mass ratio
$S=\pi\bar{r}_{h}^{2}\!+\!\frac{64\pi^{2}\epsilon\left((\beta\!+\!4\gamma)q^{4}\!-\!3\gamma q^{2}\bar{r}_{h}^{2}\!+\!\gamma\bar{r}_{h}^{4}\right)\ln\left(\bar{r}_{h}^{2}\mu_{*}^{2}\right)}{\bar{r}_{h}^{2}\left(\bar{r}_{h}^{2}-3q^{2}\right)}.$
@@
The Smarr relation can be derived from the first law via a scaling argument. In the present case, the effective action involves the dimensionful parameter $\epsilon$ and the characteristic energy scale $\mu_{*}$, so the Smarr relation must be extended by including their associated conjugate variables [39, 40]. Using the homogeneity property
$\eta\,M(S,q,\epsilon,\mu_{*})=M(\eta^{2}S,\eta q,\eta^{2}\epsilon,\eta^{-1}\mu_{*}),$ (14)
we obtain the Smarr relation for the quantum-corrected black hole:
$M=2TS+q\Phi+2V_{\epsilon}\,\epsilon-V_{\mu_{*}}\,\mu_{*},$ (15)
where $V_{\epsilon}$ and $V_{\mu_{*}}$ are computed from the free energy as $V_{\epsilon}=(\partial F/\partial\epsilon)_{T,q,\mu_{*}}$ and $V_{\mu_{*}}=(\partial F/\partial\mu_{*})_{T,q,\epsilon}$. One can verify that this Smarr relation is indeed satisfied by the thermodynamic quantities derived above.
We are now ready to explore the quantum gravitational effects on RN black holes, with a particular focus on how their charge-to-mass ratio changes [41]. In this context, we find that
$\left.\frac{\partial M}{\partial\epsilon}\right|_{T,q}=-\frac{\mathcal{N}}{300\pi}\,q^{4}\left(2\bar{r}_{h}^{2}-q^{2}\right)\frac{\ln\left(\bar{r}_{h}^{2}\mu_{*}^{2}\right)}{\bar{r}_{h}^{5}\left(3q^{2}-\bar{r}_{h}^{2}\right)},$ (16)
where $\mathcal{N}$ is defined below in eq. (20) and $\mathcal{N}>0$ always holds. The singularity at $\bar{r}_{h}=\sqrt{3}q$ indicates a critical point, which coincides with the boundary between thermodynamically stable and unstable RN black holes as determined by the sign of the heat capacity $C_{q}$. Correspondingly, for a thermodynamically stable black hole within the parameter range $q\leq\bar{r}_{h}<\sqrt{3}q$, the ratio $q/M$ increases (since $\partial M/\partial\epsilon<0$ at fixed $q$). In contrast, for an unstable black hole within $\bar{r}_{h}>\sqrt{3}q$, $q/M$ decreases. This reveals a correlation between the thermodynamic stability of RN black holes and the direction of the shift in their charge-to-mass ratio.
3.2 Extremal RN black hole and the WGC
Finally, we consider the special case of extremal RN black hole. By taking the extremal limit $T\rightarrow 0$, which is equivalent to $\bar{r}_{h}\rightarrow q$, we get
$M_{ext}=\bar{r}_{h}-\frac{\epsilon(n_{s}\!+\!6n_{f}\!+\!12n_{v}\!+\!274)\ln\left(\bar{r}_{h}^{2}\mu_{*}^{2}\right)}{600\pi\bar{r}_{h}},$
@@ 3.2 Extremal RN black hole and the WGC
$S_{ext}=\pi\bar{r}_{h}^{2}\!-\!\frac{\epsilon}{180}(n_{s}\!+\!11n_{f}\!+\!62n_{v}\!+\!424)\ln\left(\bar{r}_{h}^{2}\mu_{*}^{2}\right).$
In these equations, we have substituted the values of the Wilson coefficients $\beta$, and $\gamma$ from eqs.(4) and (5). As a consistency check, eq. (18) correctly reproduces the logarithmic correction to black hole entropy derived via an independent approach by Sen in [42], up to an additional term $3\ln A$ arising from the zero-modes associated with the rotational isometries of the extremal black hole [43].
The logarithmic entropy correction in our analysis stems from the same physical mechanism as in Sen’s approach [43], namely the one-loop effects of massless fields propagating in the black hole background. Within the effective field theory framework, these loop effects are encoded in nonlocal form factors (such as $\ln\Box$), giving rise to universal logarithmic corrections whose coefficients are determined by the low-energy particle spectrum; see also the related explicit one-loop computations for various black holes in [43, 44, 45, 46, 47].
On the other hand, logarithmic corrections to black hole entropy also emerge in other contexts with distinct physical origins. For instance, such terms arise from microscopic state counting in loop quantum gravity [48], from Cardy formula methods in conformal field theory [49], and from thermal and statistical fluctuations about the equilibrium state [50]. While these results share a common $\ln(r_{h}^{2})$ functional form, their coefficients and validity regimes are generally model- and ensemble-dependent, and do not necessarily coincide with the loop-induced coefficients derived from the low-energy effective action.
Now, from the expression of $M_{ext}$, we can straightforwardly obtain the result
@@
Intriguingly, the coefficients conspire to produce a definite sign: since $n_{s}$, $n_{f}$ and $n_{v}$ represent the numbers of particles in the spectrum, it is certain that $\mathcal{N}>0$. Consequently, the loop effects of quantum gravity surely reduce the mass of an extremal charged black hole at a fixed charge $q$, which supports the analyses presented in Refs.[12, 13]. A more detailed comparison is provided in the Appendix.
The coefficient $\mathcal{N}$ does not depend on the UV theory of quantum gravity, whereas the characteristic energy scale $\mu_{*}$ does [20]. Thus, the presence of $\ln\mu_{*}$ in eq.(19) indicates the UV sensitivity of our result. However, the model-dependent value of $\mu_{*}$ does not affect our conclusion for the WGC, since we work in the regime $r>1/\mu_{*}$ where the effective field-theoretical description is valid. Physics near the scale $r\sim 1/\mu_{*}$ would be described by a complete theory of quantum gravity and is not relevant to our current analysis.
Below we make further discussions on the computation of $\frac{\partial M_{ext}}{\partial\epsilon}$. While the above calculation is straightforward, it is worth noting that eq.(19) can also be derived from other formulas, which offer certain advantages. Concretely, we may utilize a crucial formula, proposed and proven in [51], given by
@@
In fact, from the extended first law $dM=TdS+\Phi dq+V_{\epsilon}d\epsilon$, it is straightforward to see that both sides of eq.(21) equal $\lim\limits_{T\rightarrow 0}V_{\epsilon}$. Then, from the relation $dF=-SdT+\Phi dq+V_{\epsilon}d\epsilon$, evaluating $\lim\limits_{T\rightarrow 0}V_{\epsilon}$ amounts to computing $\lim\limits_{T\rightarrow 0}\left(\frac{\partial F}{\partial\epsilon}\right)_{T,q}$, where $F=F_{RN}(T,q)-\epsilon\int d^{3}x\sqrt{g}L_{\text{hd}}$. Therefore, $\frac{\partial M_{\text{ext}}}{\partial\epsilon}$ can be evaluated as the spatial integral $\int d^{3}x\sqrt{g}L_{\text{hd}}$ for an extremal RN black hole [9, 10]. This provides a quick way to check whether the WGC is satisfied for a given higher-derivative term $\epsilon L_{\text{hd}}$.
@@
According to the formula (21), another way towards the study of the WGC is to evaluate the shift in entropy due to switching on the quantum gravity effects. The key significance of this formula lies in its ability to link the decrease in mass to the increase in entropy. Notably, in this regard, our analysis establishes a novel connection between the required super-extremality and the logarithmic entropy corrections, i.e., both of eqs.(17) and (18) are generated by the same quantum gravity effects. A caveat should be noted here: the derivative $T\left(\frac{\partial S}{\partial\epsilon}\right)_{M,q}$ is evaluated at a fixed $M$ rather than at a fixed $T$. When $\epsilon$ is switched on while keeping $M$ constant, the temperature will be shifted away from its extremal value. For this reason, we should use eq.(12) (applicable to general RN black holes) rather than the more familiar eq.(18) (which applies exclusively to extremal cases) for the calculation. It requires to use chain rule to execute the derivative and arrive at $T\left(\frac{\partial S}{\partial\epsilon}\right)_{M,q}>0$. This explains why previously no clues regarding the WGC have been derived from the logarithmic entropy corrections, despite such entropy corrections being known in the literature for many years.
4 Concluding remarks
In this paper, we have investigated the effective action obtained from integrating out massless particles at the one-loop level. Our results demonstrate that the loop effects of quantum gravity can naturally induce super-extremality ($q>M$) in charged black holes. This suggests that quantum gravity plays a crucial role in facilitating the decay of extremal black holes into more fundamental components. These effects thus have far-reaching implications for our understanding of the ultimate evolution of black holes and the universe. We emphasize that the results are model-independent, as they depend solely on the low-energy particle spectrum of nature and are insensitive to the UV details of quantum gravity theories.
Note that our new findings are mainly encapsulated in eqs.(11)–(18). Concretely, eqs.(11)–(15) provide the complete thermodynamic behaviors of the RN black hole with one-loop corrections. Regarding eq.(16), it establishes a relationship between the shift in the charge-to-mass ratio and the thermodynamic stability properties of a general RN black hole [^3] For context, a discussion on the correlation between thermodynamic stability and super-extremality can be found in Ref.[9]; however, this discussion was formalized based on tree-level higher-derivative terms for extremal black holes. Our results imply that this correlation can also be extended to one-loop effects and non-extremal cases.. Furthermore, as the extremal limit of eqs.(11) and (12), eqs.(17) and (18) reveal an unexpected connection between the super-extremality required by the WGC (reflected in the second term of $M_{ext}$) and logarithmic entropy corrections (reflected in the second term of $S_{ext}$), which were previously treated as separate research directions.
In summary, in the spirit of the swampland conjecture, any candidate for a consistent quantum gravity theory must satisfy certain fundamental physical criteria. As a pivotal principle of the swampland program, the WGC distinguishes consistent quantum gravity theories from those in the swampland. As noted in Ref.[41], the nonlocal features of loop processes remain far from well understood in the context of gravitational phenomena. Our work clarifies how the inclusion of loop effects establishes an intrinsic connection between quantum-corrected black hole thermodynamics and the WGC. Specifically, loop effects give rise to an entropy increase for a black hole with fixed mass $M$ and charge $q$. This behavior is physically intuitive: loop effects can blur the observational precision of the system’s detailed properties [9, 52]. Then, according to eq.(21), this entropy increase manifests as a corresponding mass increase. This analysis thus demonstrates that the WGC plays a fundamental role by being inherently embedded in the underlying operational framework of quantum gravity.
Acknowledgments
We thank Grant Remmen, Nima Arkani-Hamed, and Yu-Tin Huang for their constructive discussions regarding loop effects of massless particles on WGC, which took place after our work was posted on arXiv. XY would like to thank Yu Tian and Hongbao Zhang for useful discussions. YX is also thankful to the Higgs Centre for Theoretical Physics at the University of Edinburgh for providing research facilities and hospitality during the visit. This work was supported in part by the National Natural Science Foundation of China with Grant No. 12475048, the Hebei Natural Science Foundation with Grant No. A2024201012, and the China Scholarship Council with Grant No. 202408130101.
Appendix: A comparison of effective actions
Now we compare the effective action employed in our study with those adopted in previous works [12, 13]. First and foremost, the key distinction lies in the fact that our effective action is suited to consistently analyzing one-loop corrections to general RN black holes, whereas the actions in [12, 13] are constructed exclusively for the extremal case. In this Appendix, we demonstrate that our effective action is indispensable even for the extremal regime, as it is a prerequisite for establishing a self-consistent thermodynamic analysis.
@@ Acknowledgments
where $C_{\mu\nu\alpha\beta}$ represents the Weyl tensor, and the second term exhibits a structure similar to the Gauss-Bonnet term. Given that the RN black hole is Ricci-flat, two independent bases of the higher-derivative terms are sufficient in this context. Using these bases, the combinations of coefficients $\lambda_{1}=\frac{1}{2}(\beta+4\gamma)$ and $\lambda_{2}=-\frac{1}{2}(\beta+2\gamma)$ yield the values
@@
Interestingly, according to eqs.(18) and (20), we can observe that $\lambda_{1}$ governs the behavior of super-extremality, while $\lambda_{2}$ controls the logarithmic corrections to the entropy of an extremal RN black hole.
By comparison, the higher-derivative part of the effective action in Refs.[12, 13] is often expressed in terms of four-photon interactions $F^{4}$. However, we note that, through an appropriate field redefinition, it can also be transformed into the form
@@
Apparently, after our manipulation of basis choices and field redefinition, the two effective actions (22) and (25) appear nearly identical. This is not surprising since they ultimately describe the same physical phenomenon of loop effects. However, the main difference is that, in action (25), the coefficient $c_{2}\sim\lambda_{2}\ln(\bar{r}_{h}^{2}\Lambda^{2})$ explicitly contains a black hole parameter $\bar{r}_{h}$ and is simply treated as a constant in that formalism. Thus, it is only applicable for extracting information from a fixed black hole metric, i.e., the extremal RN case in the present context. However, for any issues involving variations away from the given black hole metric (for example, a self-consistent analysis of the first law of thermodynamics linking neighboring black hole solutions), this approach can easily give rise to conceptual problems when interpreting the results.
@@
Firstly, the second term of eq.(25) are often viewed as a total-derivative term and completely overlooked [13, 53]. If doing this, one can still obtain the required super-extremality of WGC (controlled by $\lambda_{1}$) but would not find the other side of the story, namely, the logarithmic corrections to the black hole entropy (controlled by $\lambda_{2}$). Secondly, one might explain the absence of $\lambda_{2}$ in eq.(17) by arguing that the second term of eq.(25) is topological and does not affect the equations of motion [12]. However, since the entropy contains a term $\lambda_{2}\ln(\bar{r}_{h}^{2}\mu_{*}^{2})$ for both the non-extremal and extremal cases, the fact is that it is essential for the black hole mass to have a corresponding term $\sim\lambda_{2}\frac{1}{\bar{r}_{h}}$ to maintain the balance (otherwise, the thermodynamic law $dM=TdS+\Phi dq$ cannot hold). This term $\lambda_{2}\frac{1}{\bar{r}_{h}}$ is not in a logarithmic form, so we did not write it explicitly in eqs.(11) and (17), but actually there it is when examining the complete formula. So one can’t simply omit the second term of eq.(25) by arguing it is topological. This indicates that the analysis should be more meticulous than initially anticipated.
@@
We now provide a consistent explanation from the perspective of our effective action (22) for why $\lambda_{2}$ appears in the logarithmic term of black hole entropy, while it doesn’t appear in the logarithmic term of the mass. The second term of eq.(22) can be decomposed into two parts: $\lambda_{2}\ln{\frac{\Box}{\mu_{*}^{2}}}=\lambda_{2}\ln{\Box}-\lambda_{2}\ln{(\mu_{*}^{2})}$. Then, in eq.(22), the terms proportional to $\lambda_{2}\ln(\mu_{*}^{2})$ indeed form the Gauss-Bonnet term. It is well-known that a Gauss-Bonnet term can cause a constant shift in the entropy, so it results in an entropy correction of the form $\sim\lambda_{2}\ln(\mu_{*}^{2})$. Due to a matching mechanism between $\ln(\bar{r}_{h}^{2})$ and $\ln(\mu_{*}^{2})$, the entropy actually receives a correction of the form $\sim\lambda_{2}\ln(\bar{r}_{h}^{2}\mu_{*}^{2})$, where $\ln(\bar{r}_{h}^{2})$ stems from the contribution of $\ln\square$. Regarding the expression of mass, the terms proportional to $\lambda_{2}\ln(\mu_{*}^{2})$ are topological and thus do not contribute to its corrections. Then the matching mechanism also prevents the presence of the term $\lambda_{2}\ln{(r_{h}^{2})}$. This explains why $\lambda_{2}$ doesn’t present in the logarithmic term of the mass. However, the terms associated with $\lambda_{2}\ln\Box$ is not truly a Gauss-Bonnet term, nothing can preclude them from modifying the metric so that it causes the mass to obtain a non-logarithmic correction of the form $\sim\lambda_{2}\frac{1}{\bar{r}_{h}}$. Such corrections happen to ensure that the thermodynamic law $dM=TdS+\Phi dq$ holds. See Ref.[23] for a detailed analysis of the variation $\delta\ln\Box$ and its impact on the field equations and black hole solutions.